OpenStudy (justmeandme):

A compound containing sodium, chlorine, and oxygen is 25.42% sodium by mass. A 3.25 g sample gives 4.33 × 10^22 atoms of oxygen.What is the empirical formula?

OpenStudy (photon336):

I was thinking we could use the fact that 1 mole of any substance has 6.02*10^23 particles and we set up a ratio to find the number of moles of Oxygen \[\frac{ 1~mole }{ 6.02*10^{23}~atoms } =\frac{ x~moles}{ 4.33*10^{22} } => 0.072~moles~O_{2}*(\frac{ 16~grams }{ mol }) = 1.15~g~O_{2}\]

OpenStudy (photon336):

as you can see, I multiplied the number of moles of oxygen by the molar mass of oxygen to get the number of grams. we know that 25.52% sodium right is in the sample so you can easily find the % of oxygen and sodium in your sample

OpenStudy (justmeandme):

How did you get 0.072 moles?

OpenStudy (photon336):

\[\frac{ 1.15 }{ 3.25 } = 35.38\] 35.38% oxygen ~ 0.3538 oxygen 25.42% sodium ~ 0.2542 sodium 39.19% chlorine ~0.3919 chlorine then we look up the molar mass of each and pretend that the % are a certain number of grams. oxygen 1/16*(35.38g) = 2.2 mol oxygen chlorine 1/35*(39.19g) = 1.11 mol Cl sodium 1/23*(25.42g) = 1.11 mol Na then we divide each one by the smallest number of moles. in this case cl (2.2)/(1.11) = 2 atoms of oxygen (1.11)/(1.11) = 1 atoms of Na (1.11/(1.11) = 1 atom of Cl \[ClO_{2}Na\]

OpenStudy (photon336):

@justmeandme I used avogadro's number to find it out 1 mole of any substance has /6.02*10^23 particles. then I set up a ratio to find the number of moles of oxygen in our sample.

OpenStudy (justmeandme):

@Photon336 would it be wrong if I chose Na to be the smallest number of moles?

OpenStudy (photon336):

nope can chose either one